OpenStudy (anonymous):

hey guyz dis question is only for who appeared iit 2012 sq root 1+x^3

OpenStudy (anonymous):



OpenStudy (anonymous):

aad r u der???

OpenStudy (anonymous):

did you???

OpenStudy (anonymous):

Square both sides.

OpenStudy (anonymous):

So, x^3 + 1 = 0 Or, (x+1) (x^2 - x+ 1) = 0

OpenStudy (anonymous):

So, x = -1

OpenStudy (anonymous):

@sheena101 , @annas , is my answer correct?

OpenStudy (apoorvk):

this is an integration right??

OpenStudy (anonymous):

But can I not solve by general method? Can anyone explain this?

OpenStudy (apoorvk):

aah.. no. integration is not that.

OpenStudy (anonymous):

Okay.

OpenStudy (anonymous):

who all appeared for IIT 2012???

OpenStudy (anonymous):

aad what r u doing?

OpenStudy (aravindg):

its very easy i didnt appear but i can help

OpenStudy (anonymous):

hahah ok ...u can try

OpenStudy (anonymous):

G r u trying?????

OpenStudy (anonymous):

apoorvk tr u der??????????

OpenStudy (apoorvk):

\[\int\limits_{}^{} (t^{-4/3} + t^{-1/3})^{1/2}.dt\] expand the term using binomial expansion for fractional expressions. and then integrate. earlier problem about that x^2/(1+x^5) mein bhi last step mein wahi hoga